OpenStudy (anonymous):

The position s (in feet) at time t (in seconds) of an object moving in a straight line is given by the equation s=2t²-5t how many seconds will it take the object to move 25 feet?

OpenStudy (mathmate):

Assuming question is asking change in distance from t=0. then solve for t in \[25=2t^2-5t\] and reject negative root.

OpenStudy (anonymous):

we are at that point then using quadratic formula.

OpenStudy (mathmate):

Yes, use quadratic formula or factorize.

OpenStudy (anonymous):

3.52?

OpenStudy (anonymous):

or 6.02

OpenStudy (mathmate):

The equation to solve is \[ 2t2−5t-25=0 \] It should give an positive integer as one of the roots.

OpenStudy (anonymous):

so we got it wrong?

OpenStudy (mathmate):

Did you factorize, or use the quad formula? If you used the quad formula, watch the sign of what's inside the square-root sign.

OpenStudy (anonymous):

we used quad formula

OpenStudy (anonymous):

let me type it

OpenStudy (anonymous):

\[\frac{ 5\pm \sqrt{-5^{2}-4(2)(-25)} }{ 2(2) }\]

OpenStudy (mathmate):

Ouch! It's \[ (-5)^2 \] instead of \[-5^2\]

OpenStudy (anonymous):

oops i meant to put that in there

OpenStudy (anonymous):

scanning what he did

OpenStudy (mathmate):

Yes, if he did (-5)^2, there would be 225 in the square-root sign, and the square-root would be 15 instead of 13.2... and the integer root follows.

OpenStudy (anonymous):

ok so 5 or 2.5

OpenStudy (anonymous):

errr -2.5 so 5

OpenStudy (mathmate):

Yep!

OpenStudy (anonymous):

thank you! posted another problem he can not even figure out how to write the equation

OpenStudy (mathmate):

You're welcome